# High-energy Emission from Turbulent Electron-ion Coronae of Accreting Black Holes

### v1

@@ Abstract
We develop a model of particle energization and emission from strongly turbulent black-hole coronae. Our local model is based on a set of 2D radiative particle-in-cell simulations with an electron-ion plasma composition, injection and diffusive escape of photons and charged particles, and self-consistent Compton scattering. We show that a radiatively compact turbulent corona generates extended nonthermal ion distributions, while producing X-ray spectra consistent with observations. As an example, we demonstrate excellent agreement with observed X-ray spectra of NGC 4151. The predicted emission spectra feature an MeV tail, which can be studied with future MeV-band instruments. The MeV tail is shaped by nonthermal electrons accelerated at turbulent current sheets. We also find that the corona regulates itself into a two-temperature state, with ions much hotter than electrons. The ions carry away roughly 60% to 70% of the dissipated power, and their energization is driven by a combination of shocks and reconnecting current sheets, embedded into the turbulent flow.
@@
software: Tristan-MP v2 (Hakobyan et al., 2023)
show]daniel.groselj@kuleuven.be
@@ I Introduction
The same process which powers the electron energization, and in turn the X-ray emission, determines also the energetics of the coronal ions (i.e., protons). The composition of black-hole coronae is presently not well known; it could be either leptonic and dominated by pairs or of the hadronic type with a substantial amount of protons. In a number of sources, the compactness $\ell$ may not be high enough to sustain a pair-plasma-dominated state (Hinkle & Mushotzky, 2021). Models focusing on electron-ion compositions are particularly timely, given the recent neutrino observations by IceCube (Abbasi et al., 2025), which suggest that AGN coronae may be efficient proton accelerators. How the released energy is partitioned between the escaping photons and cosmic-ray protons can be investigated in particle-in-cell (PIC) simulations with explicit radiative transfer, which track the self-consistent interactions among photons, charged particles and their electromagnetic fields (Grošelj et al., 2024; Nättilä, 2024). However, existing studies focus mainly on pair plasma compositions.
@@ II.1 Regime of Turbulence
The global shape and multiphase structure of black-hole coronae, including the field line geometry, is presently the subject of ongoing investigations (e.g., Jiang et al., 2019; Chashkina et al., 2021; Liska et al., 2022, 2023; Scepi et al., 2024; Sridhar et al., 2025; Krawczynski & Hu, 2025). On the other hand, from the point of view of a local model, important constraints on the regime of turbulence can be obtained based on two aspects: i) the observationally inferred large radiative compactness $\ell\gg 1$, and ii) the presumed efficient particle acceleration. Alongside recent IceCube observations (Abbasi et al., 2025), hinting at proton acceleration, measurements of an extended MeV emission tail in the stellar-mass binary Cygnus X-1 (McConnell et al., 2002; Zdziarski et al., 2017) imply that black-hole coronae can also accelerate electrons into nonthermal power laws.
@@
where $\sigma_{\rm i}=B_{0}^{2}/(4\pi n_{0}m_{\rm i}c^{2})$ is the ion magnetization. Thus, the coronal ions are moderately magnetized whereas the electrons are strongly magnetized, since $\sigma_{\rm e}=\sigma_{\rm i}m_{\rm i}/m_{\rm e}\gg 1$. Interestingly, a recent analysis of global magneto-hydrodynamic (MHD) accretion simulations found $\sigma_{\rm i}\sim 0.1$ in regions identified as the corona (Hankla et al., 2025), which is compatible with our estimate (2). We also note that the collisional coupling between ions and electrons, which is neglected in our model, can become relevant on the low side of the inferred $\sigma_{\rm i}$ range (see Sec. II.4 and Appendix B).
@@ II.2 Ion Energization
Above the thermal peak of the distribution (see Eq. (7)), the ion energy spectrum may exhibit a nonthermal tail of the form ${\rm d}N/{\rm d}E\propto E^{-p}$, where $p$ is the power-law index. In the energy range above the thermal peak but below TeV band, a hard nonthermal proton slope (with $1\lesssim p\lesssim 2$) is needed to match the observed neutrino luminosities (Murase et al., 2020; Fiorillo et al., 2024; Mbarek et al., 2024; Lemoine & Rieger, 2025; Saurenhaus et al., 2025), since it is this range that sets the available amount of power for the neutrino-producing TeV protons. Predicting neutrino emission from kinetic plasma simulations requires extrapolations to domains much larger than those accessible to present-day computing resources, which introduces a major uncertainty. In our simulations (Sec. IV), we find that ions develop broad nonthermal spectra, reaching the system-size (Hillas) limit of the simulation box, but the obtained power-law slopes are relatively steep with $p\gtrsim 3$. Macroscopically large domains, particularly those with turbulence driving on scales smaller than the system size, may present a more favorable situation, since in this case the typical ion escape time can become much longer than their acceleration time. However, with larger amounts of energy accumulated in high-energy particles, the properties of the turbulent flow could be modified via the cosmic-ray feedback (Lemoine et al., 2024). Similarly as for the thermal particles, we may therefore expect that some form of self-regulation occurs also for the cosmic-ray component (e.g., as suggested by Lemoine & Rieger, 2025). In addition, nonthermal ion acceleration can be also facilitated by shocks (Inoue et al., 2019). In particular, acceleration by multiple colliding shocks (e.g., Bykov et al., 2013) within a large turbulent volume could provide a possible mechanism for producing a hard ion power law. A more detailed investigation of these aspects requires dedicated simulations with larger boxes, which are left for the future.
@@ II.3 Comptonized Emission Spectrum
In our simulations (Sec. IV) we find that electron acceleration is enabled by collisionless reconnection at intense current sheets, which are embedded intermittently into the turbulent flow (Comisso & Sironi, 2019). Under the expected coronal conditions, reconnection of intense current sheets proceeds in the trans-relativistic and moderate-guide-field regime with $\sigma_{\rm e}\gg 1$, $\sigma_{\rm i}\sim 0.1$, and $\delta B\sim B_{0}$ (Rowan et al., 2017; Werner et al., 2018; Ball et al., 2018; Comisso, 2024). [^5] Existing intermittency models of magnetized plasma turbulence imply that the most intense structures of various sizes roughly preserve the driving-scale fluctuation amplitude (e.g., Chandran et al., 2015). Thus, when $\delta B/B_{0}\sim 1$ at the driving scale, it is reasonable to assume that the most intense current sheets feature a moderate guide field. Note that while the filling fraction of intense sheets may be small, the amount of plasma reprocessed by such sheets during their lifetime can amount to a sizable fraction of the volume (Comisso & Sironi, 2019). For these parameters, Comisso (2024) reports a broken electron power law with a hard exponent $p\approx 1$ between the thermal (i.e., Maxwellian) peak of the distribution and the break at $E_{\rm break}\simeq 0.1\sigma_{\rm e}m_{\rm e}c^{2}$, which is followed by a steeper power law with $p\approx 3.5$ up to some cutoff energy $E_{\rm max}$. An upper bound on $E_{\rm max}$ is provided by radiative cooling. By balancing the radiation drag with acceleration by the reconnecting electric field we find for $\tau_{\rm T}\simeq 1$ that (e.g., see Werner et al., 2019; Sironi & Beloborodov, 2020):
$E_{\rm max}\lesssim E_{\rm rad}\simeq m_{\rm e}c^{2}\left(\frac{\beta_{\rm rec}\sigma_{\rm e}^{1/2}s_{\rm esc}/d_{\rm e}}{f_{\rm KN}\ell}\right)^{1/2},$ (11)
@@
In a turbulent box, the expected average nonthermal spectrum is a composite between the one injected at current sheets and the one from electrons in the bulk of the turbulent volume. Outside the current sheets electrons are energized less efficiently, and so their box-averaged spectrum will be generally softer than the one produced at current sheets alone. In particular, it is reasonable to expect that the hard power-law range (below $E_{\rm break}$) may not show up clearly in the average spectrum.
@@ II.5 Alternative Dissipation Scenarios
Alongside turbulence, dissipation in the corona can be mediated by shocks and/or reconnecting current sheets. We stress that our model does not exclude these alternative dissipation mechanisms, and in fact it combines elements of both shocks and reconnecting sheets, which can form naturally within the turbulent flow. A reconnection model was developed by Beloborodov (2017) and explored in a series of numerical simulations (Sironi & Beloborodov, 2020; Sridhar et al., 2021, 2023, 2025). While current sheets can form naturally as part of the turbulent cascade, our present model differs from the work by Beloborodov (2017) in that we consider a more volume-filling dissipation scenario, rather than an isolated reconnecting current sheet in a quiescent background plasma. The latter would imply that dissipation takes place only in a small volume fraction of the corona, which would modify our estimates. Furthermore, the dynamic timescale for the turbulence model is $\sim s_{\rm esc}/v_{\rm A}$, while in the model by Beloborodov (2017) the dynamic timescale is longer by a factor of $\sim 1/\beta_{\rm rec}\sim 10$.
@@ III Numerical Method
For computational convenience, we perform the simulations in a 2D periodic box of size $S\times S$. A mean magnetic field $\mathbf{B}_{0}=B_{0}\hat{\mathbf{z}}$ is imposed in the out-of-plane $z$-direction. Turbulence is driven by applying an external time-varying electric current in the form of a “Langevin antenna” (TenBarge et al., 2014). The antenna is configured to drive large-amplitude magnetic fluctuations with $\delta B\sim B_{0}$ and polarization perpendicular to $\mathbf{B}_{0}$. We drive the external current at modes with wavenumbers $(k_{x}S/2\pi,\,k_{y}S/2\pi)=(1,0),\,(0,1),\,(1,1),\,(1,-1)$ and frequencies $\omega_{0}=\pm 1.3v_{\rm A}(2\pi/S)$. The antenna decorrelation rate (TenBarge et al., 2014) is $\gamma_{0}=0.5\omega_{0}$.
@@ IV.1 Properties of the Turbulent Flow
In Fig. 3 we analyze the 1D turbulent energy spectra ${\mathcal{E}}(k_{\perp})$ as a function of the wavenumber $k_{\perp}$ perpendicular to $\mathbf{B}_{0}$. Over the MHD scale range ($k_{\perp}\lesssim 1/d_{\rm i}$) the turbulence spectrum shown in the bottom panel of Fig. 3 forms a power law with a slope close to 3/2-3/2. This is consistent with refined models of MHD turbulence that incorporate the dynamic alignment effect (Boldyrev, 2006). Near the transition into the kinetic range the spectrum then steepens as a result of dispersive wave physics and collisionless damping of the turbulent cascade (Howes, 2015). The position of the spectral break is more closely aligned with the scale of the ion thermal gyroradius $\rho_{\rm i}=\sqrt{\beta_{\rm i}}\,d_{\rm i}$ (estimated from the steady-state ion beta $\beta_{\rm i}\approx 0.57$), rather than with $d_{\rm i}$, in accordance with theoretical expectations for low-frequency Alfvénic turbulence (Schekochihin et al., 2009). The position of the break could be also affected by shocks, which have a typical width of a few ion skin depths (Appendix A). From the steady-state spectrum we also estimate the turbulence correlation length, which can be taken as a proxy for the driving scale $s_{0}$. We find $s_{0}\simeq\pi\left(\int k_{\perp}^{\;\!-1}{\mathcal{E}}(k_{\perp}){\rm d}k_{\perp}\right)/\left(\int{\mathcal{E}}(k_{\perp}){\rm d}k_{\perp}\right)\approx S/3$, and so the effective ratio $s_{\rm esc}/s_{0}\approx 3/2$ is slightly larger than unity for our choice of the forcing.
Figure 3: Turbulent energy spectra in our simulation with $\sigma_{\rm i}=0.19$, averaged over the steady state from $t=4\,S/v_{\rm A}$ until the end of the run at $t=10.3\,S/v_{\rm A}$. In the bottom panel we show the 1D energy spectrum ${\mathcal{E}}(k_{\perp})$ as a function of the perpendicular wavenumber $k_{\perp}$. Power-law slopes are shown for reference. In the top panel we show the relative energy content of fluctuations identified as the MHD Alfvén, slow, and fast mode (see Appendix C for details).
@@
It is known that the cascade of turbulent fast modes is isotropic in full 3D geometry (Cho & Lazarian, 2003; Takamoto & Lazarian, 2016), in contrast to the Alfvénic cascade which is anisotropic in wavenumber space (Goldreich & Sridhar, 1995). Therefore, 3D kinetic simulations are required for a complete picture of how the partitioning among different modes shapes the nature of the turbulent cascade and the resulting particle energization. An exact 3D equivalent of our 2D PIC simulations is computationally infeasible at present. 3D aspects of the problem could be instead studied with reduced-kinetic or fluid-like models, provided that such approximations are appropriate for the specific question at hand.

### v2

@@ Abstract
We develop a model of particle energization and emission from strongly turbulent black-hole coronae. Our local model is based on a set of 2D radiative particle-in-cell simulations with an electron-ion plasma composition, injection and diffusive escape of photons and charged particles, and self-consistent Compton scattering. We show that a radiatively compact turbulent corona generates extended nonthermal ion distributions, while producing X-ray spectra consistent with observations. As an example, we demonstrate excellent agreement with observed X-ray spectra of NGC 4151. The predicted emission spectra feature an MeV tail, which can be studied with future MeV-band instruments. The MeV tail is shaped by nonthermal electrons accelerated at turbulent current sheets. We also find that the corona regulates itself into a two-temperature state, with ions much hotter than electrons. The ions carry away roughly two-thirds of the dissipated power, and their energization is driven by a combination of shocks and reconnecting current sheets, embedded into the turbulent flow.
@@
software: Tristan-MP v2 (Hakobyan et al., 2023), xspec (Arnaud, 1996)
show]daniel.grosel@gmail.com
@@ I Introduction
The same process which powers the electron energization, and in turn the X-ray emission, determines also the energetics of the coronal ions (i.e., protons). The composition of black-hole coronae is presently not well known; it could be either leptonic and dominated by pairs or of the hadronic type with a substantial amount of protons. In a number of sources, the compactness $\ell$ may not be high enough to sustain a pair-plasma-dominated state (Haardt et al., 1997; Hinkle & Mushotzky, 2021). Models focusing on electron-ion compositions are particularly timely, given the recent neutrino observations by IceCube (Abbasi et al., 2025), which suggest that AGN coronae may be efficient proton accelerators. How the released energy is partitioned between the escaping photons and cosmic-ray protons can be investigated in particle-in-cell (PIC) simulations with explicit radiative transfer, which track the self-consistent interactions among photons, charged particles and their electromagnetic fields (Grošelj et al., 2024; Nättilä, 2024). However, existing studies focus mainly on pair plasma compositions.
@@ II.1 Regime of Turbulence
The global shape and multiphase structure of black-hole coronae, including the field line geometry, is presently the subject of ongoing investigations (e.g., Jiang et al., 2019; Chashkina et al., 2021; Liska et al., 2022, 2023; Scepi et al., 2024; Sridhar et al., 2025; Krawczynski & Hu, 2025; Nagele et al., 2026). On the other hand, from the point of view of a local model, important constraints on the regime of turbulence can be obtained based on two aspects: i) the observationally inferred large radiative compactness $\ell\gg 1$, and ii) the presumed efficient particle acceleration. Alongside recent IceCube observations (Abbasi et al., 2025), hinting at proton acceleration, measurements of an extended MeV emission tail in the stellar-mass binary Cygnus X-1 (McConnell et al., 2002; Zdziarski et al., 2017) imply that black-hole coronae can also accelerate electrons into nonthermal power laws.
@@
where $\sigma_{\rm i}=B_{0}^{2}/(4\pi n_{0}m_{\rm i}c^{2})$ is the ion magnetization. Thus, the coronal ions are moderately magnetized whereas the electrons are strongly magnetized, since $\sigma_{\rm e}=\sigma_{\rm i}m_{\rm i}/m_{\rm e}\gg 1$. Interestingly, a recent analysis of global magneto-hydrodynamic (MHD) accretion simulations found $\sigma_{\rm i}\sim 0.1$ in regions identified as the corona (Hankla et al., 2026), which is compatible with our estimate (2). We also note that the collisional coupling between ions and electrons, which is neglected in our model, can become relevant on the low side of the inferred $\sigma_{\rm i}$ range (see Sec. II.4 and Appendix B).
@@ II.2 Ion Energization
Above the thermal peak of the distribution (see Eq. (7)), the ion energy spectrum may exhibit a nonthermal tail of the form ${\rm d}N/{\rm d}E\propto E^{-p}$, where $p$ is the power-law index. In the energy range above the thermal peak but below the TeV band, a hard nonthermal proton slope (with $1\lesssim p\lesssim 2$) is needed to match the observed neutrino luminosities (Murase et al., 2020; Fiorillo et al., 2024; Mbarek et al., 2024; Lemoine & Rieger, 2025; Saurenhaus et al., 2026), since it is this range that sets the available amount of power for the neutrino-producing TeV protons. Predicting neutrino emission from kinetic plasma simulations requires extrapolations to domains much larger than those accessible to present-day computing resources, which introduces a major uncertainty. In our simulations (Sec. IV), we find that ions develop broad nonthermal spectra, reaching the system-size (Hillas) limit of the simulation box, but the obtained power-law slopes are relatively steep with $p\gtrsim 3$. Macroscopically large domains, particularly those with turbulence driving on scales smaller than the system size, may present a more favorable situation, since in this case the typical ion escape time can become much longer than their acceleration time. However, with larger amounts of energy accumulated in high-energy particles, the properties of the turbulent flow could be modified via the cosmic-ray feedback (Lemoine et al., 2024). Similarly as for the thermal particles, we may therefore expect that some form of self-regulation occurs also for the cosmic-ray component (e.g., as suggested by Lemoine & Rieger, 2025). In addition, nonthermal ion acceleration can be also facilitated by shocks (Inoue et al., 2019; Nhat Ly et al., 2026). In particular, acceleration by multiple colliding shocks (e.g., Bykov et al., 2013) within a large turbulent volume could provide a possible mechanism for producing a hard ion power law. A more detailed investigation of these aspects requires dedicated simulations with larger boxes, which are left for the future.
@@ II.3 Comptonized Emission Spectrum
In our simulations (Sec. IV) we find that electron acceleration is enabled by collisionless reconnection at intense current sheets, which are embedded intermittently into the turbulent flow (Comisso & Sironi, 2019). Under the expected coronal conditions, reconnection of intense current sheets proceeds in the trans-relativistic and moderate-guide-field regime with $\sigma_{\rm e}\gg 1$, $\sigma_{\rm i}\sim 0.1$, and $\delta B\sim B_{0}$ (Rowan et al., 2017; Werner et al., 2018; Ball et al., 2018; Comisso, 2024). [^5] Existing intermittency models of magnetized plasma turbulence imply that the most intense structures of various sizes roughly preserve the driving-scale fluctuation amplitude (e.g., Chandran et al., 2015). Thus, when $\delta B/B_{0}\sim 1$ at the driving scale, it is reasonable to assume that the most intense current sheets feature a moderate guide field. Note that while the filling fraction of intense sheets may be small, the amount of plasma reprocessed by such sheets during their lifetime can amount to a sizable fraction of the volume (Comisso & Sironi, 2019). For these parameters, Comisso (2024) reports a broken electron power law with a hard exponent $p\approx 1$ between the thermal (i.e., Maxwellian) peak of the distribution and the break at $E_{\rm break}\simeq 0.1\sigma_{\rm e}m_{\rm e}c^{2}$, which is followed by a steeper power law with $p\approx 3.5$ up to some cutoff energy $E_{\rm max}$. An upper bound on $E_{\rm max}$ is provided by radiative cooling. By balancing the radiation drag with acceleration by the reconnecting electric field we find for $\tau_{\rm T}\sim 1$ that (e.g., see Werner et al., 2019; Sironi & Beloborodov, 2020):
$E_{\rm max}\lesssim E_{\rm rad}\simeq m_{\rm e}c^{2}\!\left[\frac{\beta_{\rm rec}(v_{\rm A}/c)\,\sigma_{\rm e}^{1/2}s_{\rm esc}/d_{\rm e}}{f_{\rm KN}\ell}\right]^{1/2},$ (11)
@@
In a turbulent box, the expected average nonthermal spectrum is a composite between the one injected at current sheets and the one from electrons in the bulk of the turbulent volume. Outside the current sheets electrons are energized less efficiently, and so their box-averaged spectrum will be generally softer than the one produced at current sheets alone. In particular, it is reasonable to expect that the hard power-law range (below $E_{\rm break}$) may not show up clearly in the average spectrum. For additional results and discussions on this matter see Appendix A.
@@ II.5 Alternative Dissipation Scenarios
Alongside turbulence, dissipation in the corona can be mediated by shocks and/or reconnecting current sheets. We stress that our model does not exclude these alternative dissipation mechanisms, and in fact it combines elements of both shocks and reconnecting sheets. A reconnection model was developed by Beloborodov (2017) and explored in a series of numerical simulations (Sironi & Beloborodov, 2020; Sridhar et al., 2021, 2023, 2025). While current sheets can form naturally as part of the turbulent cascade, our present model differs from the work by Beloborodov (2017) in that we consider a more volume-filling dissipation scenario, rather than an isolated reconnecting current sheet in a quiescent background plasma. The latter would imply that dissipation takes place only in a small volume fraction of the corona, which would modify our estimates. Furthermore, the dynamic timescale for the turbulence model is $\sim s_{\rm esc}/v_{\rm A}$, while in the model by Beloborodov (2017) the dynamic timescale is longer by a factor of $\sim 1/\beta_{\rm rec}\sim 10$.
@@ III Numerical Method
For computational convenience, we perform the simulations in a 2D periodic box of size $S\times S$. While the spatial domain is 2D, the particle velocities and electromagnetic fields have components in all 3 directions (i.e., the simulations are 2D3V). A mean magnetic field $\mathbf{B}_{0}=B_{0}\hat{\mathbf{z}}$ is imposed in the out-of-plane $z$-direction. Turbulence is driven by applying an external time-varying electric current in the form of a “Langevin antenna” (TenBarge et al., 2014). The antenna is configured to drive large-amplitude magnetic fluctuations with $\delta B\sim B_{0}$ and polarization perpendicular to $\mathbf{B}_{0}$. We drive the external current at modes with wavenumbers $(k_{x}S/2\pi,\,k_{y}S/2\pi)=(1,0),\,(0,1),\,(1,1),\,(1,-1)$ and frequencies $\omega_{0}=\pm 1.3v_{\rm A}(2\pi/S)$. The antenna decorrelation rate (TenBarge et al., 2014) is $\gamma_{0}=0.5\omega_{0}$.
@@ IV.1 Properties of the Turbulent Flow
In Fig. 3 we analyze the 1D turbulent energy spectrum ${\mathcal{E}}(k_{\perp})$ as a function of the wavenumber $k_{\perp}$ perpendicular to $\mathbf{B}_{0}$. The total turbulent energy is obtained here as the sum of energy from the electromagnetic, bulk kinetic, and density fluctuations (see Appendix C for details). Over the MHD scale range ($k_{\perp}\lesssim 1/d_{\rm i}$) the turbulence spectrum shown in the bottom panel of Fig. 3 forms a power law with a slope close to 3/2-3/2. This is consistent with refined models of MHD turbulence that incorporate the dynamic alignment effect (Boldyrev, 2006). Near the transition into the kinetic range the spectrum then steepens as a result of dispersive wave physics and collisionless damping of the turbulent cascade (Howes, 2015). The position of the spectral break is more closely aligned with the scale of the ion thermal gyroradius $\rho_{\rm i}=\sqrt{\beta_{\rm i}}\,d_{\rm i}$ (estimated from the steady-state ion beta $\beta_{\rm i}\approx 0.57$), rather than with $d_{\rm i}$, in accordance with theoretical expectations for low-frequency Alfvénic turbulence (Schekochihin et al., 2009). The position of the break could be also affected by shocks, which have a typical width of a few ion skin depths (Appendix A). From the steady-state spectrum we also estimate the turbulence correlation length, which can be taken as a proxy for the driving scale $s_{0}$. We find $s_{0}\simeq\pi\left(\int k_{\perp}^{\;\!-1}{\mathcal{E}}(k_{\perp}){\rm d}k_{\perp}\right)/\left(\int{\mathcal{E}}(k_{\perp}){\rm d}k_{\perp}\right)\approx S/3$, and so the effective ratio $s_{\rm esc}/s_{0}\approx 3/2$ is slightly larger than unity for our choice of the forcing.
Figure 3: Turbulent energy spectrum in our simulation with $\sigma_{\rm i}=0.19$, averaged over the steady state from $t=4\,S/v_{\rm A}$ until the end of the run at $t=10.3\,S/v_{\rm A}$. In the bottom panel we show the 1D energy spectrum ${\mathcal{E}}(k_{\perp})$ as a function of the perpendicular wavenumber $k_{\perp}$. Power-law slopes are shown for reference. In the top panel we show the relative energy content of fluctuations identified as the MHD Alfvén, slow, and fast mode (see Appendix C for details).
@@
It is known that the cascade of turbulent fast modes is isotropic in full 3D geometry (Cho & Lazarian, 2003; Takamoto & Lazarian, 2016), in contrast to the Alfvénic cascade which is anisotropic in wavenumber space (Goldreich & Sridhar, 1995). Therefore, 3D kinetic simulations are required for a complete picture of how the partitioning among different modes shapes the nature of the turbulent cascade and the resulting particle energization. An exact 3D equivalent of our 2D PIC simulations is computationally infeasible at present. 3D aspects of the problem could be instead studied with reduced-kinetic or fluid-like models, provided that such approximations are appropriate for the specific question at hand. This direction is left for future work.